Sally has two bags of nuts; one with seventy-nine percent peanuts, and the other with twenty-seven percent peanuts. To make a six lb bag of fifty-eight percent peanuts, how much of each should Sally use?
Answer by addingup(3677)

0.70x+0.27(6-x) = 0.58(6)
0.70x+1.62-0.27x = 3.48
0.43x = 1.86
x = 4.33
You need 4.33 lbs of 70% and 6-4.33= 1.67 lbs of the 27%
John